# www.entalus.com

### v1

@@ 9 Conclusions
Fundamental concepts of system design such as interfaces, composition, refinement, and abstraction, but also mutual recursion and corresponding induction principles are defined and derived almost seamlessly from the causality-contraction correspondence together with established results in fixpoint and approximation theory. It is well beyond the scope and the ambition of this treatise, however, to develop a comprehensive and readily applicable framework for system design. For such an endeavor the stream-based calculus for distributed, concurrent, real-time systems in [10] may serve as a starting point.
In contrast to traditional denotational approaches for complete partial orders we do not need to be constantly concerned with modeling and reasoning about undefinedness. Reasoning about undefinedness seems to be an appropriate concept for modeling concrete computations with their multitude of sources of erroneous behavior. Simultaneously chasing these all potential sources of errors and their consequences, however, tends to be overwhelming in practice. Here, principled system design offers a suitable alternative by gradually introducing ”real-world” issues and features to an abstract system design or algorithm in a compositional manner [23]. In general, this kind of incremental design can be accomplished even when features are mutually dependent [44].
@@
Quantifiable fixpoint approximation is another advantage of the proposed metric-based approach to system design. These bounds provide useful anytime information to human system designers and also to mechanized verification and design engines, such as ”the 1024 element prefix of the stream solution has already been established in the current Picard iteration”. On the other hand, these approximations are neither under- nor over-approximations according to some partial order as is the case, for example, with Kleene-style fixpoint iteration.
Disclaimer. The author thanks Manfred Broy who pointed out to him the close correspondence between the logical notion of causality and metric-based contraction. Indeed these notes grew out of curiosity in investigating at least some of the immediate consequences of this profound correspondence.
Appendix 0.A Stream Ultrametric
For the prefix metric $d$ on streams the strong triangle inequality holds.
Proposition 12
$d(f,h)$ $\leq\mathrm{max}({d(f,g)},{d(g,h)})\mbox{\@.}$ (32)
for all streams $f$, $g$, $h$.
Proof
It is the case that $f(j)=g(j)$ for $j<v({f-g})$ and $g(j)=h(j)$ for $j<v({g-h})$. By transitivity, $f(j)=h(j)$ for each $j<\mathrm{min}({v({f-g})},{v({g-h)})})$. Therefore, $v({f-h})\geq\mathrm{min}({v({f-g})},{v({g-h})})$, which is equivalent to $2^{-v({f-h})}\leq\mathrm{max}({2^{-v({f-g})}},{2^{-{v({g-h)})}}})$. Thus, inequality (32) holds.
As a consequence, the isosceles triangle principle
Corollary 8
$d(f,h)$ $=\mathrm{max}({d(f,g)},{d(g,h)})\textrm{ ~{}~{}~{}when }d(f,g)\neq d(g,h)$ (33)
holds for streams $f$, $g$, $h$.
Proof
For establishing (33) suppose $d(f,g)<d(g,h)$. Then,
$d(f,h)\leq max(d(f,g),d(g,h))=d(g,h)\mbox{\@,}$
but also, since $d(g,h)$ is not $\leq d(f,g)$, $d(g,h)\leq max(d(f,g),d(f,h))=d(f,h)$, and we are done.
Among the numbers $d(f,h)$, $d(f,g)$, $d(g,h)$ the largest and second largest are equal, and therefore (33) states that each three points are vertices of an isosceles triangle.
A consequence of (33) is: if $B(f,r_{1})$ and $B(g,r_{2})$ are two closed balls in an ultrametric space, with $r_{1}<r_{2}$, then either $B(f,r_{1})\,\cap\,B(g,r_{2})=\emptyset$ or $B(f,r_{1})\subseteq B(g,r_{2})$.
Lemma 15
$({A}^{\omega},d)$ is Cauchy complete.
Proof
Let $(f_{n})_{n\in\mathbb{N}}$ be a Cauchy sequence with $f_{n}\in{A}^{\omega}$. Then, for all $k\in\mathbb{N}$ there is $N_{k}\in\mathbb{N}$ such that $|{f_{n}-f_{m}}|<|{X^{k}}|$ for all $n\geq m\geq N_{k}$. This means that $f_{n}-f_{m}\in X^{k}{A}^{\omega}$. Writing $f_{n}=\sum_{j\in\mathbb{N}}a_{n,j}X^{j}$ we get that $(a_{n,j})_{n\in\mathbb{N}}$ is constant for $n$ large enough, and
$\lim_{n\to\infty}f_{n}=\sum_{j\in\mathbb{N}}(\lim_{n\to\infty}a_{n,j})X^{j}\in{A}^{\omega}\mbox{\@.}$
Appendix 0.B Hausdorff Distance
Proposition 13
$d(f,H)$ $=\mathrm{max}({d(f,g)},{d(g,H)})$
for all $f,g\in{A}^{\omega}$ and nonempty $H\subseteq{A}^{\omega}$.
Proof
  1. 1.
    In case $d(g,H)<d(f,g)$, then there is $h\in H$ such that $d(g,h)\leq d(f,g)$, and therefore
    $d(f,H)$ $\leq d(f,h)\leq\max{(d(f,g),d(g,h)}$
    $=d(f,g)=\max{(d(f,g),d(g,H)}$
  2. 2.
    In case $d(f,g)\leq d(g,H)$, then for all $h\in H$, $d(f,g)\leq d(g,h)$, and hence
    $d(f,H)\leq d(f,h)\leq\max{(d(f,g),d(g,h)}=d(g,h)\mbox{\@.}$
    Therefore, $d(f,H)\leq\inf_{h\in H}{d(g,h)}=d(f,H)=\max{(d(f,g),d(g,H)}$.
Lemma 16
$({\mathrm{CB}({{A}^{\omega}})},\mathcal{H}_{d})$ is an ultrametric space.
Proof
Let $F,G,H\in{\mathrm{CB}({{A}^{\omega}})}$, and let $f\in F$ be arbitrary.
  1. 1.
    In case there is a $g\in G$ with $d(f,g)\leq d(g,H)$, then
    $d(f,H)$ $\leq\max{(d(f,g),d(g,H)}=d(g,H)$
    $\leq\mathcal{H}_{d}{(G,H)}\leq\max{(\mathcal{H}_{d}(F,G),\mathcal{H}_{d}(G,H))}\mbox{\@.}$
  2. 2.
    Otherwise, in case $d(g,H)<d(f,g)$ for all $g\in G$, then
    $d(f,H)\leq\max{(d(f,g),d(g,H))}=d(f,g)\mbox{\@,}$
    and therefore
    $d(f,H)\leq\inf_{g\in G}d(f,G)\leq\mathcal{H}_{d}{(F,G)}\leq\max{(\mathcal{H}_{d}{(F,G)},\mathcal{H}_{d}{(G,H)})}\mbox{\@.}$
Similarly, $d(h,F)\leq\max{(\mathcal{H}_{d}{(F,G)},\mathcal{H}_{d}{(G,H)})}$ for each $h\in H$ and therefore
$\mathcal{H}_{d}{(F,H)}$ $\leq\max{(\mathcal{H}_{d}{(F,G)},\mathcal{H}_{d}{(G,H)})}\mbox{\@.}$
Appendix 0.C Uniform Convexity
Streams equipped with |.||{.}|| . | are uniformly convex. Expressed in geometrical terms this property is simple: ”it is that the mid-point of a variable chord of the unit sphere of the space cannot approach the surface of the sphere unless the length of the chord goes to zero” [15].
Proposition 14
$({A}^{\omega},|{.}|)$, for $A$ a field, is uniformly convex.
Proof
We need to show that for every $0<\varepsilon\leq 2$ there is some $\delta>0$ such that for any two streams $f$, $g$ on the unit ball, that is $|{f}|=|{g}|=1$, the condition $|{f-g}|\geq\varepsilon$ implies that
$|{\frac{f+g}{2}}|\leq 1-\delta\mbox{\@.}$
The condition $|{f-g}|\geq\varepsilon$ implies that $[X^{0}](f-g)=0$, and therefore $[X^{0}]f=[X^{0}]g$. Moroever, $[X^{0}]f,[X^{0}]g\neq 0$, since, by assumption, both $f$ and $g$ lie on the unit ball. Now,
$[X^{0}](\frac{f+g}{2})=[X^{0}]f\neq 0$
and the norm $|{\frac{f+g}{2}}|$ is strictly less than $1$.
As a consequence of uniform convexity, the triangle inequality for linear independent vectors is strict with a uniform bound [22].
Appendix 0.D $\varepsilon$-Chainability
For given $\varepsilon>0$, a metric space $(M,d)$ is said to be $\varepsilon$-chainable if and only if for any given $a,b\in M$ there is an $\varepsilon$-chain from $a$ to $b$, that is a finite set of points $x_{0},\ldots x_{n}\in M$ with $a=x_{0}$, $b=x_{n}$, and $d(x_{i-1},x_{i})<\varepsilon$ for all $i=1,2,\ldots,n$.
Proposition 15
$({A}^{\omega},d)$ is $\varepsilon$-chainable.
Proof
Let $\varepsilon>0$ and $f,g\in{A}^{\omega}$. Let $k$ be minimum with $2^{-k}\leq\varepsilon$. In case $f,g\neq 0$ there are $n,m\in\mathbb{N}$ with ${f}\!\!\restriction_{n}=0$ and ${g}\!\!\restriction_{m}=0$. We may assume $n<m$: now patch $f$ with trailing zeros from $n$ to $n+k$. If the distance of the patched stream to $g$ is less then $\varepsilon$ we are are, and if not we repeat patching (a finite number of times) until we get close enough to $g$ In case, $f\neq 0$, $g=0$ a similar procedure shows $\varepsilon$-chainability. The case $f=g=0$ trivially holds.
Definition 13
([38], p. 480) Let $(M,d)$ be a metric space. A multivalued function $T:M\to{\mathrm{CB}({M})}$ is said to be $(\varepsilon,\lambda)$-uniformly locally contractive mapping (where $\varepsilon>0$ and $10\leq\lambda<$ 1 provided that, if $x,y\in M$ then $\mathcal{H}_{d}({T(x)},{T(y)})\leq\lambda d(x,y)$.
Theorem 0.D.1
([38], Theorem 6, p. 481) Let $(M,d)$ be a complete $\varepsilon$-chainable metric space. If $T:M\to{\mathrm{Comp}({M})}$ is an $(\varepsilon,\lambda)$-uniformly locally contractive mapping, then $T$ has a fixpoint.
Appendix 0.E Type 2 Turing Machines
Recall that a basic open set on ${A}^{\omega}$ the product topology, and therefore also of the topology induced by $d$, is of the form $U(\alpha)\,\coloneqq\,\{{f\in{A}^{\omega}}\,\mid\,{|\alpha|=n,{f}\!\!\restriction_{n}=\alpha}\}$ for $\alpha\in A^{*}$ a finite word.
Proposition 16
Stream transformers $T:{A}^{\omega}\to{B}^{\omega}$ as computed by Type 2 Turing machines are continuous (with respect to the topology induced by $d$).
Proof
(Sketch) For a basic open set $V\,\coloneqq\,U(\alpha)$, for arbitrary finite word $\alpha\in{B}^{\omega}$, we need to verify that $W\,\coloneqq\,T^{-1}(V)$ is open. Consider $f\in W$, then $T(f)\in V$. Thus on input $f$ the machine produces an output tape starting with the finite word $\alpha$. By the time it writes out these cells it has inspected at most a finite prefix of the input tape.333 Within a particular type of space-time (which our universe may or may not possess), however, there can be situations in which a form of observer-relative acceleration occurs and an observer can witness the evolution of an entire infinite world-line by some finite time point [27]. We verify that $f\in U(\alpha)\subseteq W$. It is obvious that $f\in U(\alpha)$. To prove $U(\alpha)\subseteq W$ take any $g\in U(\alpha)$ and observe that $T(f)$ and $T(g)$ agree on the first $|\alpha|$ values of the output. This implies that $T(g)\in V$ and hence $g\in W$, as required.

### v2

@@ 9. Conclusions
Fundamental concepts of system design such as interfaces, composition, refinement, and abstraction are defined and derived from the causality-contraction connection together with established results in fixpoint and approximation theory. Moreover, the correspondence between the functional and a more machine-oriented view of causal stream transformers may advantageously be used in system design to switch between these intertwined view points as needed. It is well beyond the scope and the ambition of this exposition, however, to develop a comprehensive and readily applicable framework for system design (Broy, 2023).
@@
Quantifiable fixpoint approximation is one of the main practical advantages of the proposed metric-based approach to system design. These bounds provide useful anytime information both to human system designers and to mechanized verification and design engines; such as: ”the required 1024 element prefix of the stream solution has already been established in the current Picard iteration”.
In some cases, the causality condition on stream maps may be considered to be too restrictive as it includes continuity. This condition, however, can be relaxed since the ultrametric space of streams is $\varepsilon$-chainable (for given $\varepsilon>0$); that is, there is a finite path between any two streams with intermediate ”jumps” less than $\varepsilon$. For $\varepsilon$-chainable spaces a local contraction condition suffices to obtain fixpoints for single- (Edelstein, 1961) and multivalued maps (Nadler, 1969).
The causality-contraction connection easily, but at the expense of notational overhead, extends to transformers with a mix of discrete and dense (that is, indices are real-valued) streams in heterogeneous products of streams to model hybrid systems. Moreover, probabilistic systems may be modeled by Menger’s probabilistic metric space, or any variant thereof, in which the distance between any two points is a probability distribution function (Sherwood, 1971; Kutbi et al., 2015). Together, these developments yield a comprehensive mathematical foundation for the formal construction of cyber-physical systems (Lee, 2008; Broy, 2013) and their realization on computing machinery.